For a binomial distribution, which probability is not equal to the probability of 1 success in 5 trials where the probability of success is .4?

the probability of 4 failures in 5 trials where the probability of success is .6

the probability of 1 success in 5 trials where the probability of failure is .6

the probability of 4 failures in 5 trials where the probability of failure is .6

the probability of 4 failures in 5 trials where the probability of success is .4
Answer #1

We are given x is number of successes = 1

n is number of trials = 5

and probability of success = 0.4

Therefore probability of failure = 1 - 0.4 = 0.6

number of failures = 5-1 = 4

Therefore the probability of 1 success in 5 trials where the probability of success is .4 that is P( x = 1 ) is not equal to the Probability of 4 failures in 5 trials where the probability of success is .6  

( because probability of failure would be 1- 0.6 = 0.4 )